OpenStudy (anonymous):

If f(x) = 6x – 4, evaluate f(x)/-4 for x=8 Help Fast!!! A.–2 B.–11 C.44 D.11

OpenStudy (shamim):

f(x)=6x-4 f(8)=6*8-4 f(x)=48-4 f(x)=44

OpenStudy (anonymous):

thank you

OpenStudy (shamim):

f(x)/(-4)=44/(-4)=-11

OpenStudy (anonymous):

wait which one is right the last one?

OpenStudy (shamim):

b

OpenStudy (anonymous):

ok...

OpenStudy (shamim):

medal?

OpenStudy (anonymous):

how do u give medal?

OpenStudy (shamim):

click on best answrr

OpenStudy (anonymous):

oh yea i did

OpenStudy (anonymous):

best response

OpenStudy (shamim):

ok;

OpenStudy (anonymous):

can u look at my next question

OpenStudy (anonymous):

?